Heat Capacity of a Gas at Constant Pressure

A system comprised of 7.900×101 g of H2(g) cools from 170.0°C to 35.0°C at a constant pressure of 4.5 atm.
The molar heat capacity at constant pressure for H2(g) is 28.86 J K−1 mol−1.

a) Calculate q.
b) Calculate w.

c) Calculate delta E .

d) Calculate delta H .

Note - for part a) I have tried -1.54*10^5 J, 1.54*10^5 J and -3.08*10^5 J (these are all wrong). I'm on my last try, please help!
Answer #1

GTiven 7.900x 1D The Tempesatur T = 1%0-0c T35.0 A Conptin pressure45afm onslant paure The molay hea! FOX (ondon preare proe39.140 maley Ou (39.190)x (28:86) (350-100) (39-190) (2886)-135) = 152688-159 -152688- 16 oyle -152688.16 Touley nR (T-7 alheA E = 15 2688.16 - 33,884.95 = -186,513. Joules AE -186,573 11 Joles AH AH-a AH 15 2668 16 Jooleg